OpenStudy (anonymous):

PLZ HELP!! A cyclist rides her bike at a rate of 5 meters per second. What is this rate in kilometers per hour? How many kilometers will the cyclist travel in 3 hours? Do not round your answers.

OpenStudy (anonymous):

@thekre8or

OpenStudy (anonymous):

5 meters per second = 0.005 kilometers per second so multiply 0.005 by 3600 then multiply your answer by the three and dont round

OpenStudy (anonymous):

and that should be kilometers in 3 hours

OpenStudy (anonymous):

when i solved it was 54 kilometers

OpenStudy (anonymous):

@thekre8or 0.005 * 3600 is 18 so that is the rate km per hour?

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

just do this 1k= 1000 meter so 1meter=1/1000 1h=3600s 1 s=1/3600 so 5 m/s=(5/1000)x3600 k/h

OpenStudy (anonymous):

the multiply by 3 and u should get 54

OpenStudy (xapproachesinfinity):

always try to do proportion dealing with such problem we have rate of 18 km in 1 hour 18km=1hour x=3hours so x=3x18=54 km

OpenStudy (xapproachesinfinity):

Try always to make your life easier and simple ^_^